OpenStudy (anonymous):

Help will fan and Medal !!:)) Algebra-find the value of each variables and the measure of the angles

OpenStudy (anonymous):



OpenStudy (anonymous):

In a triangle, 3 angles add up to 180 degrees. So... y1 (left y) = 180 -28 -85 = 67 degrees you also know that y1 and y2 (right y) add up to 180 degrees y2 = 180 - 67 = 113 degrees Use the 3 angle rule again add up to 180... z = 180 - 113 - 17 = 50 degrees

OpenStudy (anonymous):

@piglet9 Thank you very much for helping me,It helps alot:))